OpenStudy (oleg3321):

Choose the slope-intercept equation of the line that passes through the point (-2, 4) and is parallel to y = -3x + 6.

OpenStudy (oleg3321):

y =1/3 x +14/3 y = 3x + 10 y = -3x - 2 y = -1/3 x + 10/3

OpenStudy (oleg3321):

@driftracer305

OpenStudy (oleg3321):

@Fifciol

OpenStudy (oleg3321):

please help

OpenStudy (oleg3321):

i just need answer

OpenStudy (driftracer305):

i would definitely pick C.... because parallel lines have same slopes no matter what

OpenStudy (oleg3321):

thanks

OpenStudy (driftracer305):

sure.....